# Cybersecurity of Quantum Key Distribution Implementations

### v1

@@ Abstract
Practical implementations of Quantum Key Distribution (QKD) often deviate from the theoretical protocols, exposing the implementations to various attacks even when the underlying (ideal) protocol is proven secure. We present new analysis tools and methodologies for quantum cybersecurity, adapting the concepts of vulnerabilities, attack surfaces, and exploits from classical cybersecurity to QKD implementation attacks. We present three additional concepts, derived from the connection between classical and quantum cybersecurity: “Quantum Fuzzing”, which is the first tool for black-box vulnerability research on QKD implementations; “Reversed-Space Attacks”, which are a generic exploit method using the attack surface of imperfect receivers; and a concrete quantum-mechanical definition of “Quantum Side-Channel Attacks”, meaningfully distinguishing them from other types of attacks. Using our tools, we analyze multiple existing QKD attacks and show that the “Bright Illumination” attack could have been fully constructed even with minimal knowledge of the device implementation. This work begins to bridge the gap between current analysis methods for experimental attacks on QKD implementations and the decades-long research in the field of classical cybersecurity, improving the practical security of QKD products and enhancing their usefulness in real-world systems.
@@ 1 Introduction
We next define a general family of attacks named Reversed-Space Attacks (Section 5), which is a generic exploit method for discovered vulnerabilities. This method, initially introduced in the preliminary work [27, 28], begins by finding the reversed space of the QKD implementation, which is derived by reversing states that are measured by the QKD device “back in time”. This space is essentially the receiving party’s attack surface, because it exactly describes the quantum space affecting their measurements. Then, the method uses the reversed space to construct a successful attack (see Appendix B for a detailed example). We note that an additional important example of a reversed-space attack is the “fixed-apparatus attack”, which appeared in  [24].
In contrast to attacks that only rely on the transmitted and received quantum states, there are also attacks that utilize unmeasured side channels; such attacks are named “side-channel attacks”. Side-channel attacks on quantum devices have appeared in the literature [20, 21, 29, 30]. However, previous definitions were too general, because all attacks on QKD systems rely on physical faults. We thus give a new and meaningful definition of Quantum Side-Channel Attacks (see Section 6).
@@ 1.1 Other Related Work
Our work focuses on QKD scheme implementations using photons. This is not a limitation, since all commercial implementations and almost all experimental implementations of QKD are photon-based [12, 13, 14, 5, 15, 33, 34, 35, 36, 37].
@@
Finally, in this paper, we use classical cybersecurity terminology that has been commonplace in the last few decades; see [25] for an educational handbook and the NIST SP 800 documents [26] for an international cybersecurity standard. Both provide documentation of this terminology, as well as a broad introduction to classical cybersecurity as a whole.
@@ 2.1 The (Ideal) BB84 Protocol
In the BB84 QKD protocol [1], Alice sends to Bob $N$ qubits, each of them randomly chosen from the set of quantum states $\{|{0}\rangle,|{1}\rangle,|{+}\rangle\triangleq\frac{|{0}\rangle+|{1}\rangle}{\sqrt{2}},|{-}\rangle\triangleq\frac{|{0}\rangle-|{1}\rangle}{\sqrt{2}}\}$, and Bob measures each of them either in the computational basis $\{|{0}\rangle,|{1}\rangle\}$ or in the Hadamard basis $\{|{+}\rangle,|{-}\rangle\}$, chosen randomly. Thereafter, Alice and Bob post-process the results by using an authenticated classical channel. If Alice and Bob used matching bases, they now share a bit (unless there is some noise or eavesdropping); if they use mismatching bases, Bob’s results are random. Alice and Bob reveal their basis choices and discard the bits for which they used mismatching bases. After that, they publicly reveal a random subset of their bits in order to estimate the error rate (then discarding those exposed test bits), aborting the protocol if the error rate is too high; and they perform error correction and privacy amplification processes, obtaining a completely secret, identical final key.
@@ 3.1 Classical Cybersecurity
    Examining the Attack Surface. The first step in analyzing a system’s security is to examine the various interfaces through which an attacker could gain access. In a standard website-based blog, an attacker can target any interface normally available to all blog readers. For example, a reader can typically post a comment by sending a “post comment” request. Sometimes, there are additional available interfaces. For instance, there might be a way to send and receive files from the blog-hosting server (also known as an FTP interface), or even access the blog administrator’s interface. These interfaces may be blocked to external connections, protected by passwords, or left completely open. Other potential interfaces, such as physical access to the server (for example, plugging in an infected USB drive), are not available to an attacker outside the premises, but could be available to an internal attacker, such as a malicious employee.
@@ 3.2.1 Vulnerabilities
A common form of vulnerability in Alice’s devices is that Alice sometimes sends different states than those she intended. A well-known example of such a vulnerability can be seen in the Photon-Number Splitting (PNS) attack [16, 48] (which showed all QKD experiments done until around year 2000 to be insecure). The PNS vulnerability is based on a fault in Alice’s transmitter that sometimes sends two-photon pulses (instead of a single photon) in one of the four allowed configurations.
@@ 3.2.2 Attack Surface
Every QKD device, be it Alice’s or Bob’s, uses a specific state space for its operation. Each state space is a component of the system’s attack surface. In Alice’s device, the state space component of the attack surface is the span of the states she transmits. For example, in the ideal BB84, Alice transmits qubits. However, implementation vulnerabilities can enlarge this space: for example, if Alice’s device suffers from the Photon Number Splitting (PNS) vulnerability, her attack surface changes from the space spanned by her attack surface becomes the space spanned by $\{|{1,0}\rangle^{\!\text{\tiny F}},|{0,1}\rangle^{\!\text{\tiny F}}\}$ to the space spanned by $\{|{0,0}\rangle^{\!\text{\tiny F}},|{1,0}\rangle^{\!\text{\tiny F}},|{0,1}\rangle^{\!\text{\tiny F}},|{2,0}\rangle^{\!\text{\tiny F}},|{1,1}\rangle^{\!\text{\tiny F}},|{0,2}\rangle^{\!\text{\tiny F}}\}$.
@@
Every device interacts with its physical environment to a certain degree. Depending on how drastically a QKD device can be affected through the interaction with the environment, or how much internal information is exposed through it, the interaction with the environment can enable critical attacks on the QKD device. For example, consider the physical configuration of Alice’s or Bob’s device when they prepare to send or measure a state. If this configuration can somehow be probed by Eve, then it is part of the attack surface of the devices, and Eve can potentially use it for an attack [17]. Section 6 defines attacks based on this part of the attack surface as “Quantum Side-Channel Attacks” and includes a detailed explanation of the above example.
@@ 3.2.3 Exploits
Since all successful attacks on QKD must be based on some imperfection in the implementation, every successful attack can be partitioned into two parts: a vulnerability (or several vulnerabilities) and an exploit that utilizes it. This decomposition of all attacks to vulnerabilities and exploits highlights, just like in classical systems, that finding a vulnerability in a QKD system is not enough to attack it. To carry out a successful attack, the vulnerability must be used through a matching exploit. From this we conclude that building exploits is a field of its own that requires careful consideration; in fact, even when a vulnerability (or several ones) is already known, building an exploit is a completely separate task which can be very challenging.
@@
A common theme in exploit research for classical cybersecurity is the idea of generic exploit methods that apply to classes of vulnerabilities, and not necessarily to one specific instance of a vulnerability. Assuming that a class of vulnerabilities trigger similar behavior in a system (often called a primitive in classical cybersecurity), one can build a generic exploit method that relies on that behavior. This idea is also commonplace in existing QKD literature discussing “attack families”.
@@ 4.1 Motivation: Practical Adversaries for QKD Systems
QKD protocols aim to provide unconditional security: the generated key should be secure without making any assumption on the capabilities or knowledge of the attacker. Similarly, when examining the security of a practical implementation, a desired goal is to guarantee the security of the implementation without imposing any conditions or limitations on Eve. In particular, Eve is assumed to be all-powerful and all-knowledgeable, where her knowledge includes the structure of the implementation, and any action Alice and Bob can take as part of the protocol. Under this assumption, we must assume that Eve has knowledge of any vulnerability that exists in the devices used by Alice and Bob, which she can and will exploit in order to conduct her attack.
In contrast, classical cybersecurity considers a wide range of possible adversaries when examining system security. Some (resourceful) adversaries have intimate knowledge of the system’s inner workings, while others do not. Furthermore, even adversaries who know the inner workings of a system may not have detected all its vulnerabilities; the same may be true even for the system designers themselves. Thus, we conclude that many practical adversaries and system designers do not know a priori all the vulnerabilities in a specified system.
@@
The same applies when analyzing the security of QKD systems: system designers and attackers might not be fully aware of physical vulnerabilities in a given implementation, making vulnerability research an important tool for identifying vulnerabilities in QKD systems.
@@ 4.2 Quantum Fuzzing: Definition
The concept of Quantum Fuzzing has previously been discussed in the context of quantum software testing [50, 51]. However, here we wish to consider it in a different context: of cybersecurity and vulnerability research for QKD.
@@ 5.4 Discussion: Countermeasures to Reversed-Space Attacks
In Appendix C we revisit the Reversed-Space Attack on interferometric BB84 that was described in Appendix B, while adding the countermeasure of measuring additional time bins in Bob’s implementation. On the one hand, our analysis shows that, under the restriction of Eve to sending single-photon pulses and non-collective attacks, Bob’s given device is effective in blocking Reversed-Space Attacks. On the other hand, we demonstrate that the additional modes measured by Bob effectively create a new implementation with a different, larger reversed space. Eve can now use a broader range of attacks than before on this new implementation.
@@ 6 Quantum Side-Channel Attacks
In this section, we examine side-channel attacks, which exploit unintended physical interfaces to compromise security, in the context of QKD devices. As classical intuition fails to capture the nature of side channels in QKD devices, we introduce a new definition of “Quantum Side-Channel Attacks”. Through concrete examples, we show that Quantum Side-Channel Attacks can target a wide range of device components. Because these attacks are often overlooked in standard security analyses, our definition emphasizes the need to treat them explicitly, providing system designers with a clearer understanding and a stronger basis for defending against them.
@@ 6.1 The Classical Concept of Side-Channel Attacks
In cybersecurity, side-channel attacks are attacks on computing systems that utilize unconventional interfaces exposed by the target system. An attack can use these interfaces to extract data from the system, as well as modify the system’s behavior.
@@
6.2 Adapting Side-Channel Attacks to QKD
@@
In a transmitting device (Alice), the conventional interface is the device’s transmissions: the space of states that the device can transmit. In a measuring device (Bob), the conventional interface is the device’s measurement of incoming signals, which is modeled as the space of states that Bob measures (or that affect Bob’s measurement). Thus, we view side-channel attacks on QKD as attacks that depend on more than Alice’s sent states, or affect something other than Bob’s measured space (or the ancillas Eve can save inside her system). We arrive at the following definition:
@@
An attack is a Quantum Side-Channel Attack if:
@@
    either its input depends non-trivially on more than Alice’s sent states,
@@
    or its output non-trivially affects more than Bob’s measured space and the attacker’s private ancillas.
A full, formal definition, based on the Hilbert spaces that each device uses, can be found in Appendix E.
There have been past attempts to classify attacks on QKD implementations that do not rely on the conventional interfaces of the devices. Vakhitov, Makarov, and Hjelme [17] used the term “conventional optical eavesdropping” to describe attacks where “…Eve can get information by using loopholes in Alice’s and Bob’s optical set-up rather than by measuring the transmitted quantum states.” This term is, in fact, a special case of our definition. While forcing leakage from the optical setups of Alice and Bob certainly qualifies as a side channel, our definition does not require that we not measure the transmitted quantum states. In fact, as shown below, several Quantum Side-Channel Attacks collect data from QKD devices through side channels, or manipulate QKD devices through side channels, in order to perform more successful measurements of the transmitted quantum states. Furthermore, there are other elements besides the optical setup that can be used to launch a Quantum Side-Channel Attack, as explored below.

### v2

@@ Abstract
Practical implementations of Quantum Key Distribution (QKD) often deviate from the theoretical protocols, exposing the implementations to various attacks even when the underlying (ideal) protocol is proven secure. We present new analysis tools and methodologies for quantum cybersecurity, adapting the concepts of vulnerabilities, attack surfaces, and exploits from classical cybersecurity to QKD implementation attacks. We also present three additional concepts, derived from the connection between classical and quantum cybersecurity: “Quantum Fuzzing”, which is the first tool for black-box vulnerability research on QKD implementations; “Reversed-Space Attacks”, which are a generic exploit method using the attack surface of imperfect receivers; and concrete quantum-mechanical definitions of “Quantum Side-Channel Attacks” and “Quantum State-Channel Attacks”, meaningfully distinguishing them from each other and from other attacks. Using our tools, we analyze multiple existing QKD attacks and show that the “Bright Illumination” attack could have been found even with minimal knowledge of the device implementation. This work begins to bridge the gap between current analysis methods for experimental attacks on QKD implementations and the decades-long research in the field of classical cybersecurity, improving the practical security of QKD products and enhancing their usefulness in real-world systems.
@@ 1 Introduction
We next define a general family of attacks named Reversed-Space Attacks (Section 5), which is a generic exploit method for discovered vulnerabilities. This method, initially introduced in the preliminary work of [27, 28], begins by finding the reversed space of the QKD implementation, which is derived by reversing states that are measured by the QKD device “back in time”. This space is essentially the receiving party’s attack surface, because it exactly describes the quantum space affecting their measurements. Then, the method uses the reversed space to construct a successful attack (see Appendix B for a detailed example). We note that an additional important example of a reversed-space attack is the “fixed-apparatus attack”, which appeared in [24].
In contrast to attacks that only rely on the transmitted and received quantum states, there are also attacks that utilize unmeasured side channels; such attacks are named “side-channel attacks”. Side-channel attacks on quantum devices have appeared in the literature [20, 21, 29, 30]; however, previous definitions of this term were too general, because all attacks on QKD systems rely on physical faults. We thus give new and meaningful definitions of Quantum Side-Channel Attacks and Quantum State-Channel Attacks (see Section 6).
@@ 1.1 Other Related Work
Our work focuses on implementations of QKD schemes using photons. This is not a limitation, since all commercial implementations and almost all experimental implementations of QKD are photon-based [12, 13, 14, 5, 15, 33, 34, 35, 36, 37].
@@
Finally, in this paper, we use classical cybersecurity terminology that has been commonplace in the last few decades. See [25] for an educational handbook and the NIST Special Publication 800-160 documents [26] for cybersecurity guidelines of the US federal government: both provide documentation of that terminology, as well as a broad introduction to classical cybersecurity as a whole.
@@ 2.1 The (Ideal) BB84 Protocol
In the BB84 QKD protocol [1], whose two legitimate participants are typically named Alice and Bob, Alice sends to Bob $N$ qubits, each of them randomly chosen from the set of quantum states $\{|{0}\rangle,|{1}\rangle,|{+}\rangle\triangleq\frac{|{0}\rangle+|{1}\rangle}{\sqrt{2}},|{-}\rangle\triangleq\frac{|{0}\rangle-|{1}\rangle}{\sqrt{2}}\}$, and Bob measures each of them either in the computational basis $\{|{0}\rangle,|{1}\rangle\}$ or in the Hadamard basis $\{|{+}\rangle,|{-}\rangle\}$, chosen randomly. Thereafter, Alice and Bob post-process the results by using an authenticated classical channel. If Alice and Bob used matching bases, they now share a bit (unless there is some noise or eavesdropping); if they use mismatching bases, Bob’s results are random. Alice and Bob reveal their basis choices and discard the bits for which they used mismatching bases. After that, they publicly reveal a random subset of their bits in order to estimate the error rate (then discarding those exposed test bits), aborting the protocol if the error rate is too high; and they perform error correction and privacy amplification processes, obtaining a completely secret, identical final key.
@@ 3.1 Classical Cybersecurity
    Examining the Attack Surface. The first step in analyzing a system’s security is to examine the various interfaces through which an attacker could interact with the system. In a standard website-based blog, an attacker can target any interface normally available to all blog readers. For example, a reader can typically post a comment by sending a “post comment” request. Sometimes, there are additional available interfaces. For instance, there might be a way to send and receive files from the blog-hosting server (such as an FTP interface), or even access the blog administrator’s interface. These interfaces may be blocked to external connections, protected by passwords, or left completely open. Other potential interfaces, such as physical access to the server (for example, plugging in an infected USB drive), are not available to an attacker outside the premises, but could be available to an internal attacker, such as a malicious employee.
@@ 3.2.1 Vulnerabilities
A common form of vulnerability in Alice’s device is that Alice sometimes sends different states than those she intended. A well-known example of such a vulnerability can be seen in the Photon-Number Splitting (PNS) attack [16, 48] (which showed all QKD experiments done until around year 2000 to be insecure). The PNS vulnerability is based on a fault in Alice’s transmitter that sometimes sends two-photon pulses (instead of a single photon) in one of the four allowed configurations.
@@ 3.2.2 Attack Surface
Every QKD device, be it Alice’s or Bob’s, uses a specific state space for its operation. Each state space is a component of the system’s attack surface. In Alice’s device, the state space component of the attack surface is the span of the states she transmits. For example, in the ideal BB84, Alice transmits qubits. However, implementation vulnerabilities can enlarge this space: for example, if Alice’s device suffers from the Photon Number Splitting (PNS) vulnerability, her attack surface changes from the space spanned by $\{|{1,0}\rangle^{\!\text{\tiny F}},|{0,1}\rangle^{\!\text{\tiny F}}\}$ to the space spanned by $\{|{0,0}\rangle^{\!\text{\tiny F}},|{1,0}\rangle^{\!\text{\tiny F}},|{0,1}\rangle^{\!\text{\tiny F}},|{2,0}\rangle^{\!\text{\tiny F}},|{1,1}\rangle^{\!\text{\tiny F}},|{0,2}\rangle^{\!\text{\tiny F}}\}$.
@@
Every device interacts with its physical environment to a certain degree. Depending on how drastically a QKD device can be affected through its interaction with the environment, or how much internal information is exposed through it, that interaction can enable critical attacks on the QKD device. For example, consider the physical configuration of Alice’s or Bob’s device when they prepare to send or measure a state. If this configuration can somehow be probed by Eve, then it is part of the attack surface of the devices, and Eve can potentially use it for an attack [17]. Section 6 defines attacks based on this part of the attack surface as “Quantum Side-Channel Attacks” and includes a detailed explanation of the above example.
@@ 3.2.3 Exploits
Since all successful attacks on QKD must be based on some imperfection in the implementation, every successful attack can be partitioned into two parts: a vulnerability (or several vulnerabilities) and an exploit that utilizes it. This decomposition of all attacks to vulnerabilities and exploits highlights that, just like in classical systems, finding a vulnerability in a QKD system is not enough to attack it. To carry out a successful attack, the vulnerability must be used through a matching exploit. From this, we conclude that building exploits is a field of its own that requires careful consideration; in fact, even when a vulnerability (or several) is already known, building an exploit is a completely separate task that can be very challenging.
@@
A common theme in exploit research for classical cybersecurity is the idea of generic exploit methods that apply to classes of vulnerabilities, and not necessarily to one specific vulnerability. Assuming that a class of vulnerabilities triggers similar behavior in a system (often called a primitive in classical cybersecurity), one can build a generic exploit method that relies on that behavior. This idea is also commonplace in existing QKD literature discussing “attack families”.
@@ 4.1 Motivation: Practical Adversaries for QKD Systems
QKD protocols aim to provide unconditional security: the generated key should be secure without making any assumption on the capabilities or knowledge of the attacker. Similarly, when examining the security of a practical implementation, a desired goal is to guarantee the security of the implementation without imposing any conditions or limitations on Eve. In particular, Eve is assumed to be all-powerful and all-knowing, where her knowledge includes the structure of the implementation, and any action Alice and Bob can take as part of the protocol. Under this assumption, we must assume that Eve has knowledge of any vulnerability that exists in the devices used by Alice and Bob, which she can and will exploit in order to conduct her attack.
In contrast, classical cybersecurity considers a wide range of possible adversaries when examining system security. Some (resourceful) adversaries have intimate knowledge of the system’s inner workings, while others do not. Furthermore, even adversaries who know the inner workings of a system may not have detected all its vulnerabilities; the same may be true even for the system designers themselves. Thus, we conclude that many practical adversaries and system designers do not know all the vulnerabilities in a specified system a priori.
@@
The same applies when analyzing the security of QKD systems in practice: system designers and attackers might not be fully aware of physical vulnerabilities in a given implementation, making vulnerability research an important tool for identifying vulnerabilities in QKD systems.
@@ 4.2 Quantum Fuzzing: Definition
The concept of Quantum Fuzzing has previously been discussed in the context of quantum software testing [50, 51]. However, here we wish to consider it in a different context: the one of cybersecurity and vulnerability research for QKD.
@@ 5.4 Discussion: Countermeasures to Reversed-Space Attacks
In Appendix C, we revisit the Reversed-Space Attack on interferometric BB84 that was described in Appendix B, while adding the countermeasure of measuring additional time bins in Bob’s implementation. On the one hand, our analysis shows that, under the restriction of Eve to sending single-photon pulses and non-collective attacks, Bob’s given device is effective in blocking Reversed-Space Attacks. On the other hand, we demonstrate that the additional modes measured by Bob effectively create a new implementation with a different, larger reversed space. Eve can now use a broader range of attacks than before on this new implementation.
@@ 6 Quantum Side-Channel Attacks
In this section, we examine side-channel attacks, which exploit unintended physical interfaces to compromise security, in the context of QKD devices. As classical intuition fails to capture the nature of side channels in QKD devices, we introduce a new definition of “Quantum Side-Channel Attacks”, and their counterpart which we name “Quantum State-Channel Attacks”. Through concrete examples, we show that Quantum Side-Channel Attacks can target a wide range of device components. Because these attacks are often overlooked in standard security analyses, our definition emphasizes the need to treat them explicitly, providing system designers with a clearer understanding and a stronger basis for defending against them.
@@ 6.1 The Classical Concept of Side-Channel Attacks
In cybersecurity, side-channel attacks are attacks on computing systems that utilize unconventional interfaces exposed by the target system. An attacker can use these interfaces to extract data from the system, as well as modify the system’s behavior.
@@
6.2 Quantum Side-Channel Attacks and Quantum State-Channel Attacks
@@
In a transmitting device (Alice), the conventional interface is the device’s transmissions: the space of states that the device can transmit. In a measuring device (Bob), the conventional interface is the device’s measurement of incoming signals, which is modeled as the space of states that affect Bob’s measurements. Thus, we view side-channel attacks on QKD as attacks that depend on more than Alice’s sent states, or affect something other than the space that affects Bob’s measurements (or the ancillas Eve can save inside her system). We arrive at the following definition:
@@
An attack is a Quantum Side-Channel Attack if either of the following two conditions is true:
@@
    its input depends on more than Alice’s sent states;
@@
    its output affects more than the space that affects Bob’s measurements and the attacker’s private ancillas.
Conversely, we define the term “Quantum State-Channel Attacks” to refer to attacks on QKD systems that are not Quantum Side-Channel Attacks and do not use elements outside of Alice and Bob’s state spaces:
Definition 9.
An attack is a Quantum State-Channel Attack if the following two conditions are both true:
  1. 1.
    its input depends only on Alice’s sent states;
  2. 2.
    its output affects only the space that affects Bob’s measurements and the attacker’s private ancillas.
Note that Quantum Side-Channel Attacks can utilize resources that cannot be used in Quantum State-Channel Attacks. In addition, there can be attacks that are neither Quantum Side-Channel Attacks nor Quantum State-Channel Attacks: for example, an attack scenario where Eve installs a camera in Bob’s lab and sees him entering his basis choices (which is typically disallowed by the assumptions of quantum cryptography).
Also note that, as an edge case, it is possible for Eve to construct an attack that only relies on Alice’s space and only affects Bob’s space, but also releases some other side-effect information that does not affect Alice or Bob. Without the side-effect, this attack would be considered a Quantum State-Channel Attack, but according to our definition, the added side-effect technically makes the full attack a Quantum Side-Channel Attack. We thus consider this edge case as a “trivial” case of Quantum Side-Channel Attacks.
A full, formal definition of the two terms, based on the Hilbert spaces that each device uses, can be found in Appendix E.
There have been past attempts to classify attacks on QKD implementations that do not rely on the conventional interfaces of the devices. Vakhitov, Makarov, and Hjelme [17] used the term “conventional optical eavesdropping” to describe attacks where “Eve can get information by using loopholes in Alice’s and Bob’s optical set-up rather than by measuring the transmitted quantum states”. This term is, in fact, a special case of our definition. While forcing leakage from the optical setups of Alice and Bob certainly qualifies as a side channel, our definition does not require that we not measure the transmitted quantum states. In fact, as shown below, several Quantum Side-Channel Attacks collect data from QKD devices through side channels, or manipulate QKD devices through side channels, in order to perform more successful measurements of the transmitted quantum states. Furthermore, there are other elements besides the optical setup that can be used to launch a Quantum Side-Channel Attack, as explored below.